how many grams of iron are produced when 45,1 g of carbon reacts with excess iron (III) oxide?


1
Expert's answer
2022-04-14T12:01:04-0400

Solution:

The molar mass of carbon (C) is 12.0107 g/mol

Therefore,

Moles of C = (45.1 g C) × (1 mol C / 12.0107 g C) = 3.755 mol C


Balanced chemical equation:

Fe2O3 + 3C → 2Fe + 3CO

According to stoichiometry:

3 mol of C produce 2 mol of Fe

Thus, 3.755 mol of C produce:

(3.755 mol C) × (2 mol Fe / 3 mol C) = 2.503 mol Fe


The molar mass of iron (Fe) is 55.845 g/mol

Therefore,

Mass of Fe = (2.503 mol Fe) × (55.845 g Fe / 1 mol Fe) = 139.78 g Fe

Mass of Fe = 139.78 g


Answer: 139.78 grams of iron (Fe) are produced